OpenStudy (anonymous):

the length of a small rectangular room is 4 units more than the width, the area of the room is 32 square units. what is the dimensions of the room? help me please!!

OpenStudy (phi):

change the English into "math" length of a small rectangular room is 4 units more than the width remove most of the words: length is 4 more than width use a letter for the unknown length and the unknown width L is 4 more than W change is to equals L = 4 more than W change more than to "add" L = 4+W

OpenStudy (phi):

next, area of rectangular room = 32 what is the formula for the area of a rectangle ? see http://www.mathopenref.com/rectanglearea.html it is area = L*W in this case, we know the area is 32, so we have L*W= 32

OpenStudy (phi):

you now have 2 equations and 2 unknowns, so you can solve them L = 4+W L*W= 32 can you solve these ?

OpenStudy (anonymous):

yes i can thanks